OpenStudy (anonymous):

Steps to solve: 3X^2+9x+6=0

OpenStudy (lgbasallote):

divide each term by 3 first. it will make your life easier

OpenStudy (anonymous):

Done that - would using the quadratic formula make it easier - is there another way to do it without using quadratic formula

OpenStudy (lgbasallote):

sure why not

OpenStudy (lgbasallote):

though you may want to know it's facctorable

OpenStudy (lgbasallote):

but yeah...to be sure quadratic formula may be better

OpenStudy (mimi_x3):

\[3x^{2}+9x+6\] \[3(x^2+3x+2)\] a*b = 2 * 1 a+ b = 2+1 \[ 3(x+1)(x+2)\]

OpenStudy (lgbasallote):

you can just get rid of the 2 @Mimi_x3 lol

OpenStudy (lgbasallote):

why waste latex :p

OpenStudy (mimi_x3):

what 2?

OpenStudy (anonymous):

so if you were doing it without the quadratic formula what would be the steps to doing it ?

OpenStudy (lgbasallote):

i meant 3

OpenStudy (mimi_x3):

how do you get rid of it?

OpenStudy (lgbasallote):

like i said divide by 3 so you ended up with x^2 + 3x + 2 right?

OpenStudy (lgbasallote):

so a = 1 b = 3 c = 2

OpenStudy (mimi_x3):

lol woops i forgot about the 0

OpenStudy (lgbasallote):

\[x = \frac{-3 \pm \sqrt{(3)^2 - 4(1)(2)}}{2(1)}\] \[x = \frac{-3 \pm \sqrt{9 - 8}}{2}\] \[x = \frac{-3 \pm \sqrt 1}{2}\] \[x = \frac{-3 \pm 1}{2}\]

OpenStudy (lgbasallote):

so \[x = \frac{-3 + 1}{2} \; ;\; \frac{-3 - 1}{2}\] \[x = \frac{-2}{2} \; ;\; \frac{-4}{2}\]

OpenStudy (anonymous):

Thank you! - i understand now - i had made a mistake right up in the begining which is why i got it wrong. Thanks

OpenStudy (lgbasallote):

you're welcome :D